The product of the page numbers on two facing pages of a book is 506. Find the page numbers?

Answer by ankor@dixie-net.com(12684)

The product of the page numbers on two facing pages of a book is 506.
Find the page numbers?
;
These would be consecutive numbers; x, (x+1)
:
x(x+1) = 506
x^2 + x - 506 = 0
:
We know it is close to the square root of 506 which is: 22.and a decimal
(x-22)(x+23) = 0
x = 22; only positive pages here
:
Pages 22 and 23